OpenStudy (devante):

whats 3 cups doubled

TheSmartOne (thesmartone):

3 cups doubled is 3 * 2 cups. What is 3 * 2 + ? Hint: 3 * 2 = 3 + 3 = ? 3 * 2 = 2 + 2 + 2 = ?

OpenStudy (otakulyfe):

6 cups